OpenStudy (anonymous):

simplify the following 7^5/7^3

OpenStudy (nincompoop):

did you read the link I gave you? http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (nincompoop):

one of the laws of exponents indicate that (x^m)/(x^n) = x^(m-n) so (5^6)/(5^3) = 5^(6-3) = 5^3